# Discrete Scale-invariant Metric Learning for Efficient Collaborative Filtering

### v1

@@ Abstract.
Metric learning has attracted extensive interest for its ability to provide personalized recommendations based on the importance of observed user-item interactions. Current metric learning methods aim to push negative items away from the corresponding users and positive items by an absolute geometrical distance margin. However, items may come from imbalanced categories with different intra-class variations. Thus, the absolute distance margin may not be ideal for estimating the difference between user preferences over imbalanced items. To this end, we propose a new method, named discrete scale-invariant metric learning (DSIML), by adding binary constraints to users and items, which maps users and items into binary codes of a shared Hamming subspace to speed up the online recommendation. Specifically, we firstly propose a scale-invariant margin based on angles at the negative item points in the shared Hamming subspace. Then, we derive a scale-invariant triple hinge loss based on the margin. To capture more preference difference information, we integrate a pairwise ranking loss into the scale-invariant loss in the proposed model. Due to the difficulty of directly optimizing the mixed integer optimization problem formulated with log-sum-exp functions, we seek to optimize its variational quadratic upper bound and learn hash codes with an alternating optimization strategy. Experiments on benchmark datasets clearly show that our proposed method is superior to competitive metric learning and hashing-based baselines for recommender systems. The implementation code is available at https://github.com/AnonyFeb/dsml.
@@ 2.1. Notations
Let users and items be denoted as $\mathcal{U}=\{1,\cdots,n\}$ and =$\mathcal{I}=\{1\cdots,m\}$ respectively. The implicit feedback $\mathcal{S}$ is defined as a subset of the cartesian product of $\mathcal{U}$ and $\mathcal{I}$: $\mathcal{S}\subseteq\mathcal{U}\times\mathcal{I}$, where $s_{ui}=1$ denotes user $u$ has an interaction/rating with item $i$; otherwise, $s_{ui}=0$. The task of personalized top-$k$ recommendation is providing a personalized ranking list of $k$ items from all items in $\mathcal{I}$ for a specific user. The proposed method aims to provide personalized recommendations with metric learning in Hamming space by projecting users and items into $d$-dimension binary codes. I define positive items of user $i$ as $\mathcal{I}_{u}^{+}=\left\{i\in\mathcal{I}:\left(u,i\right)\in\mathcal{S}\right\}\subset\mathcal{I}$; and negative items as the remaining items $\mathcal{I}_{u}^{-}=\mathcal{I}\backslash\mathcal{I}_{u}^{+}\subset\mathcal{I}$. Similarly, we define positive users who rated item $i$ as $\mathcal{U}_{i}^{+}=\left\{u\in\mathcal{U}:\left(u,i\right)\in\mathcal{S}\right\}\subset\mathcal{U}$; and negative users as other users $\mathcal{U}_{i}^{-}=\mathcal{U}\backslash\mathcal{U}_{i}^{+}\subset\mathcal{U}$.
@@ 2.2. Model Formulation
where $[\cdot]_{+}=\max\{0,\cdot\}$ represents the hinge function. $\mathbf{b}_{u}$ denotes the representation of user $i$. $\mathbf{d}_{i},\mathbf{d}_{j}$ represent vectors of positive item $i$ and negative item $j$, respectively.
@@
As illustrated in Figure 2, the objective (1) of CML aims to minimize the edge $\overline{ui}$ and push the negative sample $j$ away from $u$ with a geometrical margin $m$, that enforces $||\overline{ui}||+m\leq||\overline{uj}||$, which may drag the negative item $j$ towards to the positive item $i$. To avoid this, we penalize a longer edge $\overline{ij}$ compared to the edge $\overline{ui}$. According to the cosine rule, it can be proved that the angle surrounded by the longer edge $\overline{ui}$ and $\overline{ij}$ should be the smallest one in the triangle $\triangle uij$ (Wang et al., 2017). As the summarization of three angles in the triangle equals $180^{\circ}$, so the angle $\angle uji$ should be smaller than $60^{\circ}$, i.e., $\angle uji\leq 60^{\circ}$. Since items may belong to different categories with different intra-class variations, we aim to find an angle margin $\beta$ imposed on the angle at the negative item point $j$ instead of searching for an absolute distance margin $m$.
Given user $u$, the positive item $i$, and the negative item $j$, the proposed objective is to minimize $\angle uji$. If we decrease the angle $\angle uji$, the negative $j$ may be pulled towards the positive $i$, which violates the proposed objective. So we try to find an alternative way to push the negative item $j$ away from both the user and the positive item. The intuition is pushing negative sample $j$ away from a local distribution formed by user $u$ and the positive item $i$. Approximately, we assume the local distribution is the circle area built by the diameter $\overline{ui}$. Thus when we move the user $u$ to the center $o$ and move the positive item $i$ to the item $i^{\prime}$. $||\overline{oi^{\prime}}||=||\overline{oi}||=\frac{1}{2}||\overline{ui}||$, and $\overline{oi^{\prime}}$ is orthogonal to $\overline{oj}$. Given user $u$ and positive item $i$, if we reduce the angle $\alpha$, the negative item $j$ will be pushed away from all points in the local distribution, including $u$, $i$, $o$, and $i^{\prime}$. This motivates us to define an angle margin $\beta$ on the angle $\alpha$, and the proposed objective is
@@
(4) $\tan{\alpha}=\frac{||\overline{oi^{\prime}}||}{||\overline{oj}||}\leq\tan{\beta}.$
@@
(5) $\mathcal{L}_{scale}=\big{[}||\overline{oi^{\prime}}||^{2}-\gamma^{2}||\overline{oj}||^{2}\big{]}_{+},$
where $\gamma=\tan{\beta}$, $||\overline{oi^{\prime}}||^{2}=||\overline{oi}||^{2}=\frac{1}{4}(\mathbf{b}_{u}-\mathbf{d}_{i})^{2}$. $||\overline{oj}||^{2}=||\mathbf{d}_{j}-\frac{1}{2}(\mathbf{b}_{u}+\mathbf{d}_{i})||^{2}$.
If $\alpha\leq\beta$, i.e., $||\overline{oi^{\prime}}||^{2}\leq\gamma^{2}||\overline{oj}||$, we do not update hash codes $\mathbf{b}_{u},\mathbf{d}_{i}$, and $\mathbf{d}_{j}$; otherwise, we update them by minimizing the following scale-invariant margin loss:
(6) $\mathcal{L}_{scale}=||\mathbf{b}_{u}-\mathbf{d}_{i}||^{2}-\gamma^{2}||2\mathbf{d}_{j}-(\mathbf{b}_{u}+\mathbf{d}_{i})||^{2}.$
@@
However, the hinge loss (5) is a non-smooth objective when considering more than one negative point. So we convert the hinge loss to its smooth upper bound inspired by the recent work (Sohn, 2016). Because users and items are of constant length, i.e., $||\mathbf{b}_{u}||=||\mathbf{d}_{i}||=||\mathbf{d}_{j}||=\sqrt{d}$ in Equation(5). After dropping constant terms, Equation (5) is approximated by the following smooth log-sum-up upper bound,
@@ 3. Model Optimization
where $\pi(\xi)=\frac{1}{2\xi}\big{(}\sigma(\xi)-\frac{1}{2}\big{)}$, $\sigma(x)=\frac{1}{1+e^{-x}}$ is a sigmoid function. we call $\xi$ as the variational parameter. It’s evident that the upper bound is tight at $\xi=t$. By denoting ^$\hat{\ell}(t,\xi)=\pi(\xi)(t^{2}-\xi^{2})+\frac{1}{2}(t-\xi)+\log\big{(}1+e^{\xi}\big{)}$, we derive the following upper bound of Equation (14),
@@
[…]
@@
where $\mathbf{A}_{u}=\mathbf{A}_{xu}+4\lambda d^{2}\mathbf{A}_{yu}$, $\mathbf{c}_{u}=\mathbf{c}_{xu}+4\lambda d^{2}\mathbf{c}_{yu}$. $\mathbf{A}_{xu}$ and $\mathbf{c}_{xu}$ come from the loss ^$\hat{\ell}(x_{uij},\varphi_{uij})$, and $\mathbf{A}_{yu}$ and $\mathbf{c}_{yu}$ come from the loss ^$\hat{\ell}(y_{uij},\eta_{uij})$, and
@@
where $\mathbf{A}_{i}=\mathbf{A}_{xi}+4\lambda d^{2}\mathbf{A}_{yi}$, $\mathbf{c}_{i}=\mathbf{c}_{xi}+4\lambda d^{2}\mathbf{c}_{yi}$. Similarly, $\mathbf{A}_{xu}$ and $\mathbf{c}_{xu}$ come from the loss ^$\hat{\ell}(x_{uij},\varphi_{uij})$, and $\mathbf{A}_{yu}$ and $\mathbf{c}_{yu}$ come from the loss ^$\hat{\ell}(y_{uij},\eta_{uij})$, and
@@ 14            Update $\mathbf{d}_{i}$ by sloving the objective (19) by Xpress solver;
In the training stage, we optimize the objective (14) by a mini-batch containing multiple triplets from the same users. Specifically, we divide triplets, w.r.t users. Suppose the batch size is 16, we train the proposed model on all triplets of 16 users. In training, we optimize user by user, i.e., we train the proposed model in the way of N-pair sampling (Sohn, 2016). It considers an ($N+1$)-tuplet of training examples $\{u,i,j_{1},j_{2},\dots,j_{N-1}\}$ by sampling $N-1$ negative samples for each positive feedback $s_{ij}=1$. As evaluated in NCF (He et al., 2017), the recommender system can achieve good performance when we randomly sample four to eight negative samples for each positive feedback. So in this Chapter, we sample five negative samples for each positive item to train the proposed model.

### v2

@@ Abstract.
Metric learning has attracted extensive interest for its ability to provide personalized recommendations based on the importance of observed user-item interactions. Current metric learning methods aim to push negative items away from the corresponding users and positive items by an absolute geometrical distance margin. However, items may come from imbalanced categories with different intra-class variations. Thus, the absolute distance margin may not be ideal for estimating the difference between user preferences over imbalanced items. To this end, we propose a new method, named discrete scale-invariant metric learning (DSIML), by adding binary constraints to users and items, which maps users and items into binary codes of a shared Hamming subspace to speed up the online recommendation. Specifically, we firstly propose a scale-invariant margin based on angles at the negative item points in the shared Hamming subspace. Then, we derive a scale-invariant triple hinge loss based on the margin. To capture more preference difference information, we integrate a pairwise ranking loss into the scale-invariant loss in the proposed model. Due to the difficulty of directly optimizing the mixed integer optimization problem formulated with log-sum-exp functions, we seek to optimize its variational quadratic upper bound and learn hash codes with an alternating optimization strategy. Experiments on benchmark datasets clearly show that our proposed method is superior to competitive metric learning and hashing-based baselines for recommender systems.
@@ 2.1. Notations
[…]
@@ 2.2. Model Formulation
where $\cdot]_{+}=\max\{0,\cdot\}$ represents the hinge function. $\mathbf{b}_{u}$ denotes the representation of user $i$. $\mathbf{d}_{i},\mathbf{d}_{j}$ represent vectors of positive item $i$ and negative item $j$, respectively.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3. Model Optimization
where $\pi(\xi)=\frac{1}{2\xi}\big(\sigma(\xi)-\frac{1}{2}\big)$, $\sigma(x)=\frac{1}{1+e^{-x}}$ is a sigmoid function. we call $\xi$ as the variational parameter. It’s evident that the upper bound is tight at $\xi=t$. By denoting $\hat{\ell}(t,\xi)=\pi(\xi)(t^{2}-\xi^{2})+\frac{1}{2}(t-\xi)+\log\big(1+e^{\xi}\big)$, we derive the following upper bound of Equation (14),
@@
$\text{s.t.}\ \ \mathbf{B}\in\{\pm 1\}^{n\times d},\mathbf{D}\in\{\pm 1\}^{m\times d},\Phi\in\mathbb{R}^{|\Omega|},\mathcal{H}\in\mathbb{R}^{|\Omega|}.$
@@
where $\mathbf{A}_{u}=\mathbf{A}_{xu}+4\lambda d^{2}\mathbf{A}_{yu}$, $\mathbf{c}_{u}=\mathbf{c}_{xu}+4\lambda d^{2}\mathbf{c}_{yu}$. $\mathbf{A}_{xu}$ and $\mathbf{c}_{xu}$ come from the loss $\hat{\ell}(x_{uij},\varphi_{uij})$, and $\mathbf{A}_{yu}$ and $\mathbf{c}_{yu}$ come from the loss $\hat{\ell}(y_{uij},\eta_{uij})$, and
@@
where $\mathbf{A}_{i}=\mathbf{A}_{xi}+4\lambda d^{2}\mathbf{A}_{yi}$, $\mathbf{c}_{i}=\mathbf{c}_{xi}+4\lambda d^{2}\mathbf{c}_{yi}$. Similarly, $\mathbf{A}_{xu}$ and $\mathbf{c}_{xu}$ come from the loss $\hat{\ell}(x_{uij},\varphi_{uij})$, and $\mathbf{A}_{yu}$ and $\mathbf{c}_{yu}$ come from the loss $\hat{\ell}(y_{uij},\eta_{uij})$, and
@@ 14     Update $\mathbf{d}_{i}$ by sloving the objective (19) by Xpress solver;
In the training stage, we optimize the objective (14) by a mini-batch containing multiple triplets from the same users. Specifically, we divide triplets, w.r.t users. Suppose the batch size is 16, we train the proposed model on all triplets of 16 users. In training, we optimize user by user, i.e., we train the proposed model in the way of N-pair sampling (Sohn, 2016). It considers an ($N+1$)-tuplet of training examples $\{u,i,j_{1},j_{2},\dots,j_{N-1}\}$ by sampling N1N-1 negative samples for each positive feedback $s_{ij}=1$. As evaluated in NCF (He et al., 2017), the recommender system can achieve good performance when we randomly sample four to eight negative samples for each positive feedback. So in this Chapter, we sample five negative samples for each positive item to train the proposed model.